The management of Firebolt Industries Inc. manufactures gasoline and diesel engines through two production departments, Fabrication and Assembly. Management needs accurate product cost information in order to guide product strategy. Presently, the company uses a single plantwide factory overhead rate for allocating factory overhead to the two products. However, management is considering the multiple production department factory overhead rate method. The following factory overhead was budgeted for Firebolt:

1

Fabrication Department factory overhead

$577,200.00

2

Assembly Department factory overhead

235,200.00

3

Total

$812,400.00

Direct labor hours were estimated as follows:

Fabrication Department 5,200 hours
Assembly Department 4,900
Total 10,100 hours

In addition, the direct labor hours (dlh) used to produce a unit of each product in each department were determined from engineering records, as follows:

Production Departments Gasoline Engine Diesel Engine
Fabrication Department 2.8 dlh 1.9 dlh
Assembly Department 1.9 2.8
Direct labor hours per unit 4.7 dlh

4.7 dlh

a. Determine the per-unit factory overhead allocated to the gasoline and diesel engines under the single plantwide factory overhead rate method, using direct labor hours as the activity base . If required, round all per-direct labor hours and per-unit answers to the nearest cent.

Gasoline engine per unit
Diesel engine

per unit

b. Determine the per-unit factory overhead allocated to the gasoline and diesel engines under the multiple production department factory overhead rate method, using direct labor hours as the activity base for each department. If required, round all per-unit answers to the nearest cent.

Gasoline engine per unit
Diesel engine per unit
Answer #1

Solution:

a.) Per Unit factory Overhead Allocation Using Plant wide Overhead Rate
Total Overhead Cost = $577,200 + $235,200 = $812,400
Total Direct Labor hours = $5,200 + $4,900 = 10,100
Plant wide Overhead Rate = $812,400/10,100 = $80.4

Per Unit Overhead
Gasoline Engine Overhead = $80.4 * 4.7 = $377.9 per unit
Diesel Engine Overhead = $80.4 * 4.7 = $377.9 per unit

b.) Per Unit Factory Overhead Allocation Using Multiple Production Department Overhead Rate

Particulars fabricating Department Assembly Department
Factory Overhead $577,200 $235,200
Direct Labor Hours 5,200 4,900
Overhead rate per labor hour $111 $48

Per Unit Overhead
Gasoline Engine Overhead = $111*2.8 + $48*1.9 = $402 per unit
Diesel Engine Overhead = $111*1.9 + $48*2.8 = $345.3 per unit